OpenStudy (anonymous):

Let * be a binary operation on the set Q of rational nos defined by a*b = ab/5. Write the identity for * if any .

OpenStudy (anonymous):

Basically you just have to find e such that eb=be=b for any b in Q, according to that binary relation. In this case it's 5, because 5b/5=b5/5=b.

OpenStudy (anonymous):

@nbouscal can you actually explain this sum step by step .. i dunno the basics of binary operation .. can you teach ? please ?

OpenStudy (anonymous):

Not sure how to go more step-by-step than I did. We're having a discussion on binary operations here though: http://openstudy.com/updates/501e0ca6e4b0be43870e9a57